OpenStudy (anonymous):

how do you find f"(x) for f(x) = (4x+5) ^3

OpenStudy (anonymous):

Is that all you were given?

OpenStudy (campbell_st):

this is the chain rule y = (f(x))^n y' = n(f(x))^{n -1} * f'(x) simplify then repeat for y"

OpenStudy (anonymous):

yes, you need to find the first derivative and then the second derivative

OpenStudy (campbell_st):

thats correct

OpenStudy (anonymous):

yes, but that chain rule function doesnt help me, I have a hard time understanding it without the numbers

OpenStudy (campbell_st):

ok... so look at it this way 1st derivative if you have y = x^n y' = n *x^{n -1} does that make sense...?

OpenStudy (campbell_st):

so looking at your problem \[y = (4x + 5)^3.....y' = 3 \times (4x + 5)^2\] does that make sense so far...

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

then you multiply by the derivative of 4x + 5 which is 4 so its \[y' = 3 \times(4x +5)^2 \times 4\] giving you \[y' = 12(4x + 5)^2\] does that make sense

OpenStudy (anonymous):

yes that does

OpenStudy (campbell_st):

then 2nd derivative is found in the same way, except you are multiplying by 12 so \[y" = 12 \times 2 \times(4x + 5) \times 4\] just simplify that for the answer. with the can rule, whats inside the brackets never changes. you just keep multiplying by its derivative

OpenStudy (campbell_st):

hope that helps

OpenStudy (campbell_st):

you may need to distribute for the final answer, but I wouldn't .

OpenStudy (anonymous):

thank you!